Esperantologio / Esperanto Studies 2 (2001), 53–57

Creoles or planned languages: which have the simpler grammar?

Liu Haitao reviews the doctoral dissertation of Anett Heil^{1}

W e often hear that there are many similarities between creoles and planned

languages. Several studies on this topic are available from creolists and interlinguists (Corsetti 1999, Mühlhäusler 1992, Philippe 1991). But these studies are not satisfactory for a thorough understanding of the relationship: we need more detailed work. By approaching the problem from sociolinguistics or general linguistics, one can search for a common basis of the two fields and try to assess the usefulness of interlinguistics and creolistics for general linguistics and other branches of linguistics.

The grammar of creoles and planned languages is more reduced than that of their source languages. However, do we know for sure whether a creole is grammatically more reduced than a planned language? If so, how large is the reduction? Are there reasons to make us believe the following claims?

There have been several attempts to invent artificial languages for international communication, of which Esperanto is probably the most famous and successful—but none of these have much resemblance to pidgins, and all have actually been more complex than any of the known pidgins. (Sebba 1997)

But that simplicity does not mean that the language we construct is to be a kind of “Pidgin” incapable of expressing nuances of thought which are necessary to highly cultivated Europeans. [...] the interlanguage I am advocating in this book is totally different from such languages through being expressive and efficient, though extremely simple in its grammatical structure. (Jespersen 1928)

Analysis and description alone are not enough to judge who is right in these two contrasting viewpoints. To answer that question, we first need a quantitative analysis and a comparison between creoles and planned languages.

The book under review is Anett Heil’s doctoral dissertation at Rostock University, Germany (1998). Dr. Heil tells us that both creoles and planned languages are reduced languages, but born and developed from different directions, characterized by “naturalness” and “artificiality.” By comparing them the author hopes to shed light on the development of languages as symbolic systems. Her introduction provides factual information on three creoles (those of Mauritius, Réunion, and Haiti) and three planned languages (Esperanto, Ido, and Latino sine flexione), primarily from a sociolinguistic viewpoint. It seems that the two fields are not in balance in

1Heil, Anett (1999), Grammatische Reduktion in Frankokreolsprachen und Plansprachen. Rostocker Romanistische Arbeiten. Band 2. Frankfurt/Main: Peter Lang. 220 pp. ISBN 3-631-34692-1. the author’s heart, since she devotes considerably more attention to the creoles than to the planned languages. For example, she quotes references on general properties of creoles (pp. 16–19) that would be useful to a reader wishing to understand the issues, but for planned languages such information is missing. Moreover, we are given little information about the linguistic situation of planned languages; at least Esperanto could have received a much more detailed discussion. It is not clear to the reviewer why the author has chosen to focus on these three planned languages. Esperanto and Latino sine flexione certainly represent two typical and contrasting planned languages, but for many reasons Interlingua (IALA) seems a better candidate than Ido.

In her third chapter, Mündlichkeit versus Schriftlichkeit, the author contrasts oral and written forms of languages. Generally speaking, creoles develop from oral to written form: there are non-written creoles. Planned languages, on the other hand, begin in written form, and only then proceed to oral use. It is not difficult to find planned language projects which have remained on paper.

The author goes on to explain the principles by which she selected her texts. We note that in all the languages considered, except for Latino sine flexione, she has chosen a dramatic text. To compare the reduction in grammar with the source language, a dramatic text in French is also included. We know that there is a big difference between creoles and planned languages in oral/written development. Drama is a product of the oral use of the language; in other words, this selection of texts may favor creoles, because planned languages are used chiefly in written form. Of course, drama is a relatively acceptable choice, the sentences being usually simple. The phrases in a drama tend to be shorter and structurally simpler than in most other forms of documented language; although accessible in written from, they are closer to spoken language. From each text, ten sentences are chosen as samples for morphosyntactic structural analysis.

The fifth chapter, Morphosyntax, is among the most important in the book. Here, the author compares in detail the morphosyntactic structure of her three creoles and three planned languages. The main sources for the planned languages are Blanke (Internationale Plansprachen, 1985), Janton (Einführung in die Esperantologie, 1993), Kotzin (Geschichte und Theorie des Ido, 1916), and Peano (De Latino sine flexione, 1903). The author is of the opinion that creole morphology and syntax are more reduced than in their source language, French. A possible explanation for this is that creoles originated in a contact between two languages with the aim of facilitating interlingual communication.

Theoretically, the same word may play many roles in a creole sentence, functioning as verb, noun or adjective, and rendered unambiguous through the use of prepositions or postpositions. There are no fixed terminations for the word-class and syntactic functions in creole morphology. Like French, French-based creoles have SVO word order: Subject (nominal phrase) + predicate (verbal phrase) + object (nominal phrase) + adverbial modifier.

The author cites interlinguists who see Esperanto and Ido as strongly agglutinative languages with some properties of isolating languages, a view that is supported by reference to Zamenhof’s first book on Esperanto (1887). Since Esperanto has several invariable terminations for word classes and syntactic functions, its word order is relatively free compared to creoles and Latino sine flexione. The author does not mention that Esperanto is also a marked SVO language, as we can learn, for example, from Gledhill’s corpus-based grammar (2000). She goes on to compare the order of sentence structure in creoles and planned languages, from “nominal phrase” to “adverbial modifier.” It seems that creoles have a more complex deep structure than planned languages, but at surface levels the opposite is the case. In creoles, syntactic functions are often realized by strict word order and functional words, as in isolating languages. But in planned languages, chiefly in Esperanto, these functions are fulfilled by special syntactic markers. If we use methods and theories developed for non-isolating languages to analyze creoles, we will probably not be able to fully capture their complexities. There can be no denying, however, that Esperanto has a more complete structure than creoles, in particular regarding tense, mood and aspect in verbs.

Chapter 6, Strukturanalyse, presents a detailed analysis of grammatical categories in ten sentences in seven languages (the six already mentioned, plus French), and this provides a foundation for quantitative comparison in the seventh chapter, Vergleich der Strukturanalysen, which examines the complexity of word classes, phrases and sentences in all seven languages. As already mentioned, Esperanto has more complex markers than the other languages considered (except French). For instance, Esperanto has plural and accusative endings in -j and -n, respectively, for both nouns and adjectives, thereby increasing the grammatical complexity of its nouns, whereas in creoles these functions are realized by individual words. Does this mean that creoles are simpler than Esperanto? Unfortunately, the author offers little by way of guidance on the matter. As a measure of grammatical complexity, she calculates the quotient of markers per word for seven languages. The results are as follows.

# of markers # of words  Quotient
French  293  105  2.79
Mauritius creole  131  122  1.07
Réunion creole  167  115  1.45
Haiti creole  98 or 129  121  0.81 or 1.07 (average 0.94)
Esperanto  211  104  2.01
Ido  180  124  1.45
Latino sine flexione  3  45  0.07 (first 5 sentences)
Latino sine flexione  28  50  0.56 (last 5 sentences)

The results are of course dependent on the chosen sample of ten sentences, but the conclusion is reasonable and supports the qualitative description. If the author had used the same text translated into the various languages she examines, the outcome would have been even more convincing. French holds the highest position in the list, which is perhaps not surprising. The three creoles have values lower than 2.00; of them Réunion creole has the highest value. This is noteworthy if we consider that some linguists support the idea that the Réunion language is just a variation of standard French and not a true creole.

While we praise grammatical endings for allowing for more flexible expression in Esperanto, they do increase its complexity at least on the surface. Ido scores lower than Esperanto in this regard. Latino sine flexione exhibits its qualities as a “language without grammar” (lingua sine grammatica). For Peano, grammar is a useless device: “Grammatica, tormento de pueritia, es quasi semper inutile pro intelligentio. Viatore in natione extraneo stude pauco vocabulo de isto natione, et sine grammatica construe phrasi, que publico intellege” (1910, quoted from Barandovská-Frank 1995:117). Says Jespersen, “Peano’s ideal would be no grammar, or what he thinks is the same thing, the Chinese grammar” (1928). Interlinguists, especially the creators of planned languages, fantasize about languages without grammar, but of course there is no such language, and any claim in that direction can only be taken metaphorically. For example, in Chinese (and other isolating languages) word order plays an important role in grammatical analysis. In fact, analysis of Chinese syntactic structure using methods and theories appropriate to European languages will not reveal its complexity. The application of computational linguistics to Chinese has shown that the syntactic structure of an isolating language is more difficult to process in a computer. Falk’s study of “language without grammar” (1995) is helpful in understanding the concept of reduced grammar in general.

The eighth and ninth chapters summarize the main results of the study. Heil’s conclusions suggest that Sebba is right, because Esperanto has a higher value than creoles, while Latino sine flexione has a lower value. Ultimately, the question boils down to whether isolating languages are easier to learn than languages with other typological properties. There is no easy answer. If we are to believe the conclusion of this book, Latino sine flexione is the most reduced (simple) language from a grammatical viewpoint. Why then was Latino sine flexione less successful than Esperanto, even than Ido and Interlingua (IALA)? The reader may reply by claiming that extralinguistic factors often decide the success or failure of a language as a tool of international communication, but the author provides little information on these matters.^{2} While it is essential to provide detailed comparative grammatical information on creoles and planned languages, such a simple comparison is not enough.

The bibliography covers all the important works on creolistics, but these works contain virtually no significant discussion of planned languages. Although German is one of the main languages for interlinguistics and planned languages, and several studies on creoles and planned languages have been written in planned languages (chiefly Esperanto), the author does not use or refer to them.

In conclusion then, the thesis under review offers a detailed morphosyntactic comparison between French-based creoles and three planned languages, as well as a qualitative analysis of the syntactic complexity of these languages. All these considerations provide a good foundation for further research. I recommend the book to everyone who is interested in planned languages and creoles.

2We discuss the relationship between planned languages and creoles from the point of view of sociolinguistics and general linguistics in Liu (2001).

Creoles and planned languages 57

References

Barandovská-Frank, Véra 1995 La latina kiel interlingvo / Latein als internationale Sprache. Prague: Akademia Libroservo. 193 pp.

Corsetti, Renato 1999 Kreoliĝo de Esperanto inter personaj gustoj kaj oficiala normo. In: D. Blanke; R. McCoy; O. Buller (Eds.): Por aktiva lingvopolitiko, pp. 43–60. Rotterdam: Universala Esperanto-Asocio.

Falk, Julia S. 1995 Words without grammar: Linguists and the international auxiliary language movement in the United States. Language & Communication, 15 (3), pp. 241– 259.

Gledhill, Christopher 2000 The Grammar of Esperanto. A corpus-based description. Munich: Lincom Europa. Second edition. 151 pp.

Jespersen, Otto

1928  An international language. London: Allen and Unwin. 196 pp.

Liu Haitao 2001 Pidgins, Creoles and planned languages: Language Evolution under special conditions. In: Klaus Schubert (Ed.), Planned Languages: From Concept to Reality. Brussels: Hogeschool voor Wetenschap en Kunst. Also in: Interface. Journal of Applied Linguistics. 15 [2], 2001.

Mühlhäusler, P. 1992 Shared properties of pidgins and artificial languages. In: Tom Dutton; Malcolm Ross; Darrell Tryon (red.): The language game: papers in memory of Donald C. Laycock, pp. 283–290. Pacific Linguistics, C-110.

Philippe, Benoît

1991  Kreolsprache und Plansprache. Vienna: IEM. 24 pp.

Sebba, Mark

1997  Contact Languages: Pidgins and Creoles. New York: St. Martin Press. 304 pp.

Liu Haitao is a senior engineer in information technology. He has published articles on computational linguistics and interlinguistics. He is interested in general problems of interlinguistics and its relation to other sciences.

Author’s address: Jisuanji Zhan, Qinghai Lu Chang, Xining, Qinghai, CN-810108, China E-mail: htliu@iname.com URL: http://htliu.163.net

Received 2000 03 01. Accepted 2000 03 07